C_{60}^{+} AND C_{60}^{−} IN NEON AND ARGON MATRICES

\underline{BASTIAN KERN}, DMITRY STRELNIKOV, PATRIK WEIS, ARTUR BÖTTCHER and MAN-FRED M. KAPPES, KIT Karlsruhe Institute for Technology, Division of Physical Chemistry of Microscopic Systems, Fritz-Haber-Weg 2, 76131 Karlsruhe, Germany.

Mass-selected C_{60} anions have been deposited in noble gas matrices (Ne,Ar) kept at 6K. The vibronic properties of the species (C_{60}^{+}, C_{60}^{0}, C_{60}^{−}) were determined by NIR and MID-IR absorption spectroscopy. By co-depositing electrons without the addition of electron scavengers or donors to the host matrix, we were able to vary the C_{60}^{+}/C_{60}^{−} ratio. By changing the C_{60}^{+}/C_{60}^{−} ratio, previously known vibrational data of the C_{60}^{+} and C_{60}^{−} species could be confirmed and also extended further. DFT calculations performed on isolated C_{60}^{+} and C_{60}^{−} ions support the experimental assignment.